# ODoQ: Oblivious DNS-over-QUIC

### v1

@@ Abstract
Index Terms:
DNS, security, privacy, oblivious, encryption
@@ I Introduction
Previously, DNS used plain text queries to convert domain names to IP addresses which posed security threats and put users’ privacy at risk due to the lack of encryption on DNS data. The plain text DNS data made it easy for attackers to intercept and track the websites users visit leading to vulnerabilities such as cache poisoning attacks. In these attacks, an attacker injects false records into an RRs cache for a few target domains, thereby redirecting users to phishing websites when queried. Attackers might use the DNS traffic to understand user’s behavior by tracking the websites they visit. To overcome these security challenges and safeguard the DNS data from eavesdroppers, several protocols have been proposed and standardized by the Internet Engineering Task Force (IETF) [^1] https://www.ietf.org/. These security measures consist of protocols such, as DNS Security Extension (DNSSEC) [1], DNS-over-TLS (DoT) [2], DNS-over-DTLS (DoDTLS) [3], DNS-over-HTTPS (DoH) [4], Oblivious DNS (ODNS) [5], Oblivious DoH (ODoH) [6], and DNS-over-QUIC (DoQ) [7]. To enhance privacy, RRs use QNAME minimization [8] to only disclose parts of domain names during resolution to nameservers, as illustrated in Figure 1.
@@
The IETF has standardized DoT, DoH, and DoQ protocols as secure alternatives for plain text DNS. These protocols secure communication between the client and the RR. However, the RR continues to possess information about the client’s identity, which creates possible privacy concerns. If the RR gets compromised, it might utilize the client’s identity to monitor browsing behavior or even misuse this data for target advertisements. To enhance user privacy, Oblivious DNS (ODNS)[5] and Oblivious DoH (ODoH)[6] have been proposed, which use an intermediary proxy server to hide the client’s identity from getting exposed to the RR. The client uses the RR’s public key to encrypt the DNS query ensuring that it gets decrypted only by the RR. The proxy knows only the IP address of the client and does not know the domain it is querying, as the proxy cannot decrypt it. Whereas, the RR knows only the domain name, but not the IP address of the client, as the proxy removes the client’s IP address and transmits the packet containing the encrypted DNS request to the RR for resolution. The proxy and RR must be managed by different organizations to avoid the correlation of DNS messages, which could lead to the identification of the client.
@@
A recent study [9] found that DoQ is roughly $10\%$ faster than DoH, and despite adding some encryption overhead, DoQ is merely $2\%$ slower compared to plain text DNS over UDP. This renders DoQ a very effective and trustworthy protocol for sending DNS messages between the client and the RR. Nonetheless, a significant drawback of DoQ is that it fails to preserve the client’s identity from the RR. This privacy concern drives us to suggest an enhancement to the DoQ protocol that includes a proxy server as an intermediary, which would preserve the client’s anonymity by concealing the client’s identity from the RR.
@@ II Related Work
Figure 2 charts the DNS evolution. Several protocols have been proposed over the past decade to secure the DNS data, and to preserve the client’s information (such as the IP address and the domain name) during the client-RR communication. These protocols that secure the DNS data between the client and the RR include DoT [2], DoH [4], DoDTLS [3] (an experimental protocol that uses UDP for transport, unlike DoT which uses TCP) and DoQ. On the other hand, ODNS [5] and ODoH [6] protocols focus on preserving the client’s privacy from the RR by introducing an intermediary proxy server.
@@
DNS-over-TLS (DoT) [2] was standardized by the IETF in $2015$ under RFC $7858$. In this protocol, the client establishes a TCP three-way handshake with the DNS RR, followed by a TLS handshake. DoT messages are typically transmitted over port $853$, although either party may agree to use a different port if preferred. The chosen port must vary from port $53$ to avoid conflicts, minimizing DNS interference and lowering the risk of downgrade attacks. Given that DoT employs a comparable message format to standard plain text DNS over UDP, it is quite easy to implement as a secure alternative.
@@
DNS-over-HTTPS (DoH) [4] was standardized by the IETF in $2017$ under RFC $8484$. In this protocol, the client first establishes a three-way TCP handshake with the DNS RR, and then encapsulates the DNS query within an HTTP packet, which is transmitted to the RR. DoH offers a degree of DNS query encryption comparable to the DoT, but utilizes port $443$ for its communication. The benefit of utilizing port $443$ is that an eavesdropper cannot differentiate between a standard HTTPS packet and a DoH packet.
@@
Both DoT and DoH protect privacy by employing the proven encryption methods of TLS, although this results in extra delays when resolving DNS queries. The implementation of TCP contributes to latency (due of its three-way handshake) and head-of-line blocking (since it requires packet ordering). Furthermore, TLS adds its own processing and setup time for DNS resolution. The DoQ [7] protocol tackles these overheads by utilizing QUIC (Quick UDP Internet Connections), a transport layer protocol based on UDP created by Google. Originally, gQUIC was created to enhance video streaming performance and was subsequently accepted and formalized by the IETF as RFC 9000 [10]. Afterwards, the QUIC protocol was modified to accommodate various protocols, such as DNS, and was formalized by the IETF as RFC 9001 [11], using TLS v1.3. Additionally, QUIC enables the multiplexing of packets over various streams within a single connection, thereby eliminating the head-of-line blocking problem found in TCP. DoQ [7] leverages the QUIC protocol to enhance the security and privacy of DNS packets by placing them inside QUIC packets.
@@
As the use of encrypted DNS and public DNS resolvers to improve privacy increases, a significant issue emerges: although clients’ DNS information is secured from unauthorized access, the RR still retains access to the clients’ identities. This prompts one to consider whether it is wiser to rely on an external provider—typically situated in another country with a distinct legal framework—than on the local Internet Service Provider (ISP) of the client. To prevent the RR from acquiring full knowledge of the client’s IP address and domain name, protocols such as ODNS [5] and ODoH [6] utilize a proxy server that serves as an intermediary between the client and the RR to preserve client’s identity from getting revealed to the RR.
@@ III Proposed Approach:
    Client initiates DNS query using Hybrid Public Key Encryption (HPKE [12])
@@
    • The connection can be over a TLS 1.3 handshake using encrypted Server Name Indication (ESNI) [13] or Encrypted Client Hello (ECH) [14].
A recent study [15] performed a comprehensive study to assess the performance, robustness and deployment of the DoQ protocol in the wild. The study findings reveal that the DoQ adoption in the RRs is still restricted and only AdGuard [^2] https://adguard-dns.io/kb/general/dns-providers/ is one of the first public providers to adopt and provide DoQ functionality. These findings help implement our ODoQ protocol and observe its performance with the existing DoT, DoH, DoQ, ODNS and ODoH protocols. We believe that the two-fold strategy of our ODoQ protocol – preserving client’s privacy and encrypted communication over QUIC – increases the privacy and security of DNS data.
@@ IV Conclusion and Future Directions
In this paper, we proposed the Oblivious DNS-over-QUIC (ODoQ) protocol, utilizing the benefits of the QUIC protocol in conjunction with an intermediary proxy server to safeguard client privacy from the recursive resolver (RR). The paper describes the secure message exchange among the client, proxy, and RR, in which DNS data is encrypted both ways —- client to RR and from RR to client —- utilizing the RR’s public key and the client’s symmetric key correspondingly. We also address measures to manage decryption failure at the RR in situations where key pairs have been updated. As a future direction, we intend to create a prototype of the ODoQ protocol utilizing existing tools and public resolvers that facilitate DoQ. In addition, we aim to study and compare the overhead and performance of ODoQ with the existing DNS over encryption protocols.

### v2

@@ I Introduction
Previously, DNS used plain text queries to convert domain names to IP addresses which posed security threats and put users’ privacy at risk due to the lack of encryption on DNS data. The plain text DNS data made it easy for attackers to intercept and track the websites users visit leading to vulnerabilities such as cache poisoning attacks. In these attacks, an attacker injects false records into an RRs cache for a few target domains, thereby redirecting users to phishing websites when queried. Attackers might use the DNS traffic to understand user’s behavior by tracking the websites they visit. To overcome these security challenges and safeguard the DNS data from eavesdroppers, several protocols have been proposed and standardized by the Internet Engineering Task Force (IETF) [^1] https://www.ietf.org/. These security measures consist of protocols such, as DNS Security Extension (DNSSEC) [DNSSEC_RFC], DNS-over-TLS (DoT) [DoT_RFC], DNS-over-DTLS (DoDTLS) [DoDTLS_RFC], DNS-over-HTTPS (DoH) [DoH_RFC], Oblivious DNS (ODNS) [schmitt2018oblivious], Oblivious DoH (ODoH) [ODoH_RFC], and DNS-over-QUIC (DoQ) [DoQ_RFC]. To enhance privacy, RRs use QNAME minimization [QNAME_RFC] to only disclose parts of domain names during resolution to nameservers, as illustrated in Figure 1.
@@
The IETF has standardized DoT, DoH, and DoQ protocols as secure alternatives for plain text DNS. These protocols secure communication between the client and the RR. However, the RR continues to possess information about the client’s identity, which creates possible privacy concerns. If the RR gets compromised, it might utilize the client’s identity to monitor browsing behavior or even misuse this data for target advertisements. To enhance user privacy, Oblivious DNS (ODNS)[schmitt2018oblivious] and Oblivious DoH (ODoH)[ODoH_RFC] have been proposed, which use an intermediary proxy server to hide the client’s identity from getting exposed to the RR. The client uses the RR’s public key to encrypt the DNS query ensuring that it gets decrypted only by the RR. The proxy knows only the IP address of the client and does not know the domain it is querying, as the proxy cannot decrypt it. Whereas, the RR knows only the domain name, but not the IP address of the client, as the proxy removes the client’s IP address and transmits the packet containing the encrypted DNS request to the RR for resolution. The proxy and RR must be managed by different organizations to avoid the correlation of DNS messages, which could lead to the identification of the client.
@@
A recent study [kosek2022dns] found that DoQ is roughly $10\%$ faster than DoH, and despite adding some encryption overhead, DoQ is merely $2\%$ slower compared to plain text DNS over UDP. This renders DoQ a very effective and trustworthy protocol for sending DNS messages between the client and the RR. Nonetheless, a significant drawback of DoQ is that it fails to preserve the client’s identity from the RR. This privacy concern drives us to suggest an enhancement to the DoQ protocol that includes a proxy server as an intermediary, which would preserve the client’s anonymity by concealing the client’s identity from the RR.
@@ II Related Work
Figure 2 charts the DNS evolution. Several protocols have been proposed over the past decade to secure the DNS data, and to preserve the client’s information (such as the IP address and the domain name) during the client-RR communication. These protocols that secure the DNS data between the client and the RR include DoT [DoT_RFC], DoH [DoH_RFC], DoDTLS [DoDTLS_RFC] (an experimental protocol that uses UDP for transport, unlike DoT which uses TCP) and DoQ. On the other hand, ODNS [schmitt2018oblivious] and ODoH [ODoH_RFC] protocols focus on preserving the client’s privacy from the RR by introducing an intermediary proxy server.
@@
DNS-over-TLS (DoT) [DoT_RFC] was standardized by the IETF in $2015$ under RFC $7858$. In this protocol, the client establishes a TCP three-way handshake with the DNS RR, followed by a TLS handshake. DoT messages are typically transmitted over port $853$, although either party may agree to use a different port if preferred. The chosen port must vary from port $53$ to avoid conflicts, minimizing DNS interference and lowering the risk of downgrade attacks. Given that DoT employs a comparable message format to standard plain text DNS over UDP, it is quite easy to implement as a secure alternative.
@@
DNS-over-HTTPS (DoH) [DoH_RFC] was standardized by the IETF in $2017$ under RFC $8484$. In this protocol, the client first establishes a three-way TCP handshake with the DNS RR, and then encapsulates the DNS query within an HTTP packet, which is transmitted to the RR. DoH offers a degree of DNS query encryption comparable to the DoT, but utilizes port $443$ for its communication. The benefit of utilizing port $443$ is that an eavesdropper cannot differentiate between a standard HTTPS packet and a DoH packet.
@@
Both DoT and DoH protect privacy by employing the proven encryption methods of TLS, although this results in extra delays when resolving DNS queries. The implementation of TCP contributes to latency (due of its three-way handshake) and head-of-line blocking (since it requires packet ordering). Furthermore, TLS adds its own processing and setup time for DNS resolution. The DoQ [DoQ_RFC] protocol tackles these overheads by utilizing QUIC (Quick UDP Internet Connections), a transport layer protocol based on UDP created by Google. Originally, gQUIC was created to enhance video streaming performance and was subsequently accepted and formalized by the IETF as RFC 9000 [gQUIC_RFC]. Afterwards, the QUIC protocol was modified to accommodate various protocols, such as DNS, and was formalized by the IETF as RFC 9001 [QUIC_RFC], using TLS v1.3. Additionally, QUIC enables the multiplexing of packets over various streams within a single connection, thereby eliminating the head-of-line blocking problem found in TCP. DoQ [DoQ_RFC] leverages the QUIC protocol to enhance the security and privacy of DNS packets by placing them inside QUIC packets.
@@
As the use of encrypted DNS and public DNS resolvers to improve privacy increases, a significant issue emerges: although clients’ DNS information is secured from unauthorized access, the RR still retains access to the clients’ identities. This prompts one to consider whether it is wiser to rely on an external provider—typically situated in another country with a distinct legal framework—than on the local Internet Service Provider (ISP) of the client. To prevent the RR from acquiring full knowledge of the client’s IP address and domain name, protocols such as ODNS [schmitt2018oblivious] and ODoH [ODoH_RFC] utilize a proxy server that serves as an intermediary between the client and the RR to preserve client’s identity from getting revealed to the RR.
@@ III Proposed Approach:
    Client initiates DNS query using Hybrid Public Key Encryption (HPKE [HPKE_RFC])
@@
    • The connection can be over a TLS 1.3 handshake using encrypted Server Name Indication (ESNI) [ESNI_RFC] or Encrypted Client Hello (ECH) [ECH_RFC].
A recent study [kosek2022one] performed a comprehensive study to assess the performance, robustness and deployment of the DoQ protocol in the wild. The study findings reveal that the DoQ adoption in the RRs is still restricted and only AdGuard [^2] https://adguard-dns.io/kb/general/dns-providers/ is one of the first public providers to adopt and provide DoQ functionality. These findings help implement our ODoQ protocol and observe its performance with the existing DoT, DoH, DoQ, ODNS and ODoH protocols. We believe that the two-fold strategy of our ODoQ protocol – preserving client’s privacy and encrypted communication over QUIC – increases the privacy and security of DNS data.
@@
(a) TCP + TLS 1.2
(b) TCP + TLS 1.3
(c) QUIC
Figure 4: Handshake process of (4(a)) TCP + TLS 1.2, (4(b)) TCP + TLS 1.3, and (4(c)) QUIC connection.
Figure 4 illustrates the round-trip times (RTTs) required to establish a connection between the client and the server for various protocols. For TCP + TLS 1.2 (Figure 4(a)), the process requires 1 RTT for TCP and 2 RTTs for TLS 1.2. TCP + TLS 1.3 (Figure 4(b)) reduces this to 1 RTT for TCP and 1 RTT for TLS 1.3. Finally, QUIC (Figure 4(c)) requires 1 RTT for QUIC protocol. DoT and DoH require 3 RTTs using TLS 1.2, involving one RTT for the TCP handshake and two for the TLS handshake. TLS 1.3 cuts this down to 2 RTTs, one for TCP and one for TLS. DoQ has a quicker process, requiring just 1 RTT for the handshake because of QUIC’s combined transport and encryption. ODoH introduces an intermediary proxy server between the client and the resolver, causing communication between client-proxy and proxy-resolver to require 2 RTTs along with an extra overhead from HTTP/2. ODoQ utilizes a proxy alongside QUIC to decrease latency to only 2 RTTs required for handshakes among the client, proxy, and resolver.
@@ IV Conclusion and Future Directions
In this paper, we proposed the Oblivious DNS-over-QUIC (ODoQ) protocol, utilizing the benefits of the QUIC protocol in conjunction with an intermediary proxy server to safeguard client privacy from the recursive resolver (RR). The paper describes the secure message exchange among the client, proxy, and RR, in which DNS data is encrypted both ways – client to RR and from RR to client – utilizing the RR’s public key and the client’s symmetric key correspondingly. We also address measures to manage decryption failure at the RR in situations where key pairs have been updated. As a future direction, we intend to create a prototype of the ODoQ protocol utilizing existing tools and public resolvers that facilitate DoQ. In addition, we aim to study and compare the overhead and performance of ODoQ with the existing DNS over encryption protocols.